### v1

@@ 3.2.2 Modeled Layers
(i) Overlap runtime
@@
Figure 5: Theoretical performance model: modeling baseline and overlap runtime from individual kernel runtime.
@@
Figure 5 illustrates the modeling approach for baseline and overlap runtime, incrementally constructed with individual kernel performance assessments.
@@
Initially, we model the performance of each individual kernel - GEMM, Attention, and RNG — calculating their theoretical runtime based on bounding each by the limiters. For workload sizes within the range of typical networks (GPT Brown (2020) and Llama Touvron et al. (2023)), we found that GEMM runtime is bounded by MMA, Attention by RF bandwidth and the Issue stage, and RNG by the Issue stage and ALU pipe.
@@
Next, we integrate additional components to derive runtime estimations for composed kernels. For the Attention kernel with the element dropping step, we derive its runtime by adding the silicon-measured dropping overhead to the standalone Attention kernel runtime (Figure 5d). We also compute the runtime of the fused Attention kernel with RNG by integrating both sets of instructions and identifying the primary limiter, which typically is the Issue Stage, with the ALU pipe and RF bandwidth as close secondary limiters (Figure 5e).
@@
We then calculate the runtime of RNG and GEMM separately while the other is running concurrently. RNG runtime is based on the stand-alone RNG, scaled by the GEMM interference overhead measured in silicon. If RNG’s runtime exceeds GEMM’s, the remaining RNG operations continue at full speed once GEMM completes (Figure 5f). Similarly, we model the GEMM runtime affected by RNG interference (Figure 5g).
@@
Finally, we determine the baseline runtime from the standalone GEMM and the fused Attention-and-RNG kernel runtime (Figure 5h). The overlap runtime is derived from the maximum runtime of GEMM and RNG with interference, as Attention depends on both GEMM and RNG outputs, plus the standalone Attention runtime with only the element dropping step (Figure 5i).
4 Results
This section presents the performance outcomes of our RNG and GEMM overlapping experiments. The analysis is based on the theoretical results, supported by silicon measurements. We base our experiments on a GPT-3-like network architecture, using a common hidden dimension per head of 128, and a batch size of 1. We vary the sequence lengths from 2048 to 65536, and number of heads from 48 to 128. These conclusions are broadly applicable across various network architectures.
We validated the theoretical model’s accuracy with a mere 2% average difference between the theoretical model and silicon measurements for overlapping QKV_GEMM with RNG. The difference is calculated by averaging the absolute difference of speedup between silicon and theoretical results. With such validation, we extend our performance analysis to overlapping RNG with all four GEMM layers between the previous and the current Attention layer. Figure 6 shows the modeled overlap speedup for different sequence length and number of heads.
Figure 6: Overlap speedup across different sequence lengths and number of heads, using our theoretical performance model.
The results demonstrate significant performance improvements, with speedups up to 1.23x across the five key layers of a Transformer Block (four GEMM and one Attention layer). Specific speedups include 1.06x for GPT-3, 1.14x for LLAMA2, and 1.13x for MoE. Further analysis reveals that speedup is closely correlated with the ratio of sequence length to the number of heads, given that the batch size and the hidden dimension don’t change. The greatest improvement occurs within a specific range (Region 2 in Figure 6); the speedup gradually decreases towards Region 1 or Region 3.
@@ 4 Results
To understand this trend, we examined the runtime dependencies of each kernel on sequence length and number of heads:
@@
(a) Individual kernel runtime with different number of heads.
(b) Individual kernel runtime with different sequence length.
Figure 7: Kernel runtime variation with different sequence length and number of heads, measured on silicon. GEMM runtime scales quadratically with number of heads, whereas Attention and RNG runtime scales quadratically with sequence length.
Figure 15 presents the runtime variation for each kernel across different sequence lengths and number of heads. The data shows that while GEMM runtime scales quadratically with the number of heads, the Attention and RNG runtime scales linearly. Conversely, Attention and RNG runtime scales quadratically with sequence length, whereas GEMM runtime scales linearly.
Figure 8: Analysis of speedup across three regions.
Analysis of the three regions highlighted in Figure 6 indicates:
@@
  • Region 2 (optimal speedup): Balances sequence length and number of heads, where RNG runtime is shorter but close to the GEMM runtime. Maximum benefits from hiding RNG latency by overlapping with GEMM.
@@
  • Region 3 (decreasing speedup): RNG runtime exceeds GEMM because of the long sequence lengths, leading to full exposure of RNG operations post-GEMM completion, diminishing the overlapping benefits.
Our analysis illustrates the conditional effectiveness of the overlapping strategy based on network parameters, offering insights for optimizing LLM training performance.
5 Discussion
In this section, we dive into the stand-alone RNG’s HBM capacity requirement, and the overlapping methodology’s implication on different RNG implementations and future hardware architecture generations.
5.1 HBM Capacity Requirement for Stand-alone RNG
An overhead brought by our methodology is storing the RNG bits generated in HBM to be later used by the Attention kernel. For each element in Attention’s intermediate matrix, assuming RNG generates 1 bit per element to indicate the dropping status, we require storage of $B*nH*SQ^{2}$ in HBM, which may initially appear to be critical.
Figure 9: HBM capacity requirements for stand-alone RNG across different network configurations.
Figure 9 illustrates the HBM requirements for stand-alone RNG when the entire network is run on a single GPU. If we hypothetically allocate 8GB for RNG data, it becomes apparent that applying this methodology to networks with sequence lengths of 32K or more is not feasible on a single GPU.
However, the typical deployment scenario for LLM training involves multiple GPUs, where the workload is divided into smaller segments distributed across the GPUs. This division significantly reduces the HBM capacity needed for RNG on each GPU. Common strategies for parallelism include Tensor Parallelism Shoeybi et al. (2019) and Sequence Parallelism Korthikanti et al. (2023), which split the head and sequence length dimensions, respectively. Figure 9 also demonstrates the reduced HBM requirements when employing these parallelism mechanisms. For the three mainstream LLM training networks, GPT3, LLAMA2 and GPT4-MoE, the required HBM capacity is decreased by a factor of ten or more, depending on the chosen parallelism strategy and dimension.
It is worth noting that parallelism does not change the performance benefit from overlapping RNG and GEMM analyzed earlier. Since the workload is split evenly on each GPU, the ratios of RNG, GEMM and Attention kernel runtime stay the same.
If parallelism across multiple GPUs is still not enough to alleviate HBM storage concerns, an additional strategy can be applied which involves pipelining the RNG, GEMM, and Attention kernels. Figure 10 illustrates this strategy. It involves processing only a portion of the total computation per kernel at a time, splitting along the sequence length dimension to avoid creating dependencies in the GEMM kernel.
Figure 10: Pipelining kernels to further reduce HBM storage requirements.
Moreover, as depicted previously in Figure 2, implementing parallelism introduces communication layers within each Transformer Block. Fortunately, the runtime impact of these communication layers can also be minimized using the same pipelining technique. Since communication layers utilize distinctly separate resources from GEMM (such as NVLink bandwidth), their concurrent execution will not introduce resource contention and will effectively hide communication latency.
5.2 Implication of Cheaper RNG
@@
In our prior discussions, we focused on the use of Philox 7 for RNG implementation. This subsection explores the implications of adopting more cost-effective RNG implementations, namely Philox 5 and Philox 3, which involve fewer computational iterations and shorter runtimes.
@@
We implemented all three Philox variants (Philox 3, 5, 7) on silicon using the GH100 with the same experimental setup as previously discussed. We show representative results collected for sequence length = 16K in Figure 11, which has a consistent trend with other configurations not shown on the graph.
@@
Figure 11: Silicon measurements of stand-alone RNG kernel runtime with different Philox implementations for a sequence length of 16K.
Our theoretical performance model was validated against the silicon results on overlapping RNG with QKV_GEMM, showing an error margin consistent with previous validations. Using the validated model, we further analyzed the implications of varying RNG complexities on overlapping RNG with all four GEMM kernels in the Transformer Block.
Figure 12: Cheaper RNG implementation results in smaller overall speedup.
As the complexity of the RNG algorithm decreases, the potential for speedup through overlapping also diminishes because the runtimes of the Attention and GEMM kernels remain unchanged. Thus, cheaper RNG implementations should result in smaller overall savings, as shown in Figure 12. This is the case in most scenarios, as shown in Figure 13. However, in certain cases (such as Philox 7 implementation with number of heads = 48 and sequence length = 16K) the RNG runtime exceeds the GEMM runtime, introducing performance loss from fully exposed RNG after GEMM completes. This typically occurs with a smaller number of heads and a relatively long sequence length - Region 3 as discussed in Figure 6.
@@
Figure 13: Overlap speedup with different RNG Philox implementations, derived from theoretical model.
5.3 Hardware Exploration
Our analysis so far has been in the context of NVIDIA’s GH100 GPUs NVIDIA (f). With our fine-grained performance model, we now explore how variations in hardware design might influence the efficiency of our overlapping strategy.
A key aspect of hardware evolution, particularly with NVIDIA GPUs, is the consistent increase of computational efficiency with each new generation. This trend continues with the newest Blackwell GPUs NVIDIA (a), which motivates us to model the potential impacts of further advancements in computation power, such as enhanced FP operations or lower FP precisions.
Figure 14: Overlap speedup with different GPU compute capability (TFLOPs). While GEMM runtime decreases, RNG and Attention runtimes remain constant, leading to a greater proportional speedup.
Figure 15b illustrates the implications of increased compute capability on our overlapping technique. We maintain the assumption that computation remains the bottleneck for GEMM, with memory performance keeping pace with MMA improvements through reduced precision or increased memory bandwidth. However, other non-Tensor related limiters like the issue pipeline and the ALU pipeline are likely to remain unchanged, meaning the most significant runtime reduction would be observed in the GEMM kernel. This shift will make RNG latency an even more critical bottleneck, proportionally increasing its impact on end-to-end network performance. Although the absolute runtime difference between baseline and overlapped configurations remains similar, the relative speedup should improve.
We evaluated the theoretical performance of our overlapping strategy using a hypothetical model of a more advanced GPU, which offers twice the compute capability of the H100. Figure 15 shows that overall speedup increases up to 10% with higher compute efficiency for a variation of sequence lengths and number of heads. Our findings indicate that while reductions in GEMM runtime boost the overall speedup ratio, this benefit is primarily observed in workloads with shorter sequence lengths. For longer sequences, where RNG and Attention dominate network runtime, overlapping a shorter GEMM could exacerbate problems by fully exposing RNG latency once GEMM computation completes. Thus, it is more advantageous to overlapping RNG and GEMM computation for hardware with high computational efficiency and workloads featuring relatively shorter (8K or smaller) sequence lengths.
(a) Overlap speedup on GH100.
(b) Overlap speedup on a hypothetical GPU where the GEMM compute capability is doubled and non-Tensor limiters remain the same.
Figure 15: Overlap speedup increases with more GEMM compute capability on relatively short sequence lengths and varying number of heads.
This analysis underscores the need for next-generation hardware to consider optimizing traditionally non-Tensor related factors. On the other hand, if hardware optimization on these factors are not possible, we call for future hardware-software co-design to value kernel overlapping: efficient libraries can be developed to facilitate the overlapping of non-GEMM operations, which typically do not consume extensive computation and memory resources, with GEMM operations.
6 Related Work

### v2

@@ 3.3. Methodologies to model kernels
(i) Fusion runtime
(j) Overlap runtime
@@
Figure 4. Theoretical performance model: modeling baseline, fusion and overlap runtime from individual kernel runtime.
@@
Figure 4 illustrates the modeling approach for baseline (sequential), fusion and overlap runtime, incrementally constructed with individual kernel performance assessments.
@@
First, we model the performance of each individual kernel - GEMM, Attention, and RNG — by calculating their theoretical runtime based on bounding each by the limiters. For workload sizes within the range of typical networks (GPT (OpenAI, 2023) and Llama (Grattafiori et al., 2024)), we found that GEMM runtime is bounded by MMA, Attention by RF bandwidth and the Issue stage, and RNG by the Issue stage and ALU pipe.
@@
Next, we integrate additional components to derive runtime estimations for composed kernels. For the Attention kernel with the element dropping step, we derive its runtime by adding the dropping overhead to the standalone Attention kernel runtime (Figure 4d). We also compute the runtime of the fused Attention kernel with RNG by integrating both sets of instructions and identifying the primary limiter, which typically is the Issue Stage, with the ALU pipe and RF bandwidth as close secondary limiters (Figure 4e).
@@
We then calculate the runtime of RNG and GEMM separately while the other is running concurrently. RNG runtime is based on the stand-alone RNG, scaled by the GEMM interference overhead. If RNG’s runtime exceeds GEMM’s, the remaining RNG operations continue at full speed once GEMM completes (Figure 4f). Similarly, we model the GEMM runtime affected by RNG interference (Figure 4g).
@@
Finally, we determine the baseline runtime from three stand-alone kernels (Figure 4h), and the fusion runtime from the stand-alone GEMM and the fused Attention-and-RNG kernel runtime (Figure 4i). The overlap runtime is derived from the maximum runtime of GEMM and RNG with interference, as Attention depends on both GEMM and RNG outputs, plus the standalone Attention runtime with only the element dropping step (Figure 4j).
In the model, the ratios of overhead and kernel interference are derived from our silicon measurements. Because kernel computation patterns are regular and the bottlenecks are stable, the overhead and interference ratios can be generalized to different implementations. We averaged the numbers across various input workload sizes and model architecture configurations to find the scaling factor that was applied to our model.
4. Results
This section presents the performance results of our RNG and GEMM overlapping experiments. We discuss insights derived from our analytical model, and its validation with a silicon implementation. While the analysis presented is based on the Transformer architecture running on H100 GPUs, these conclusions are broadly applicable across various software implementations and hardware configurations. Further analysis of implications on different hardware and software is discussed in Section 5.
In our study, we assume that GPU resources for LLM training are dedicated, since it’s impractical to run other applications in parallel to such a demanding workload. That said, the analytical model is also flexible to model concurrent workloads and shared resources by modifying the configurations to carve out resources for other applications.
4.1. Performance Analysis
We show our experiments on the Transformer network architecture, using a common head dimension of 128, and a batch size of 1. We vary the sequence lengths from 2K to 64K, and embedding dimension from 4K to 16K (number of heads from 32 to 128). We use the tile size 128x128x128 for GEMM and 64x128x128 for RNG and Attention for illustration purposes. These numbers are obtained from the production-level GEMM, Attention, and dropout kernel code, which are chosen for best GPU hardware resources utilization. The exact tile size is not a key factor of the proposal we are making in this paper, and changing the tile sizes in a reasonable range does not impact the final conclusion derived.
Figure 5. Fusion and overlap speedup compared to a baseline sequential kernel implementation for a Transformer Block, obtained from our analytical model.
The results demonstrate moderate performance improvements for state-of-the-art RNG-Attention fusion, but significant improvements for proposed RNG-GEMM overlap over the fusion methodology. Specifically, for a Llama3-like workload configuration, we observe 1.26x speedup for the overlap methodology compared to the baseline, and 1.22x on top of the fusion methodology. The performance benefits remain similar as models grow larger: for a later GPT4-like configuration, we observe 1.25x speedup for the overlap methodology compared to the baseline, which is 1.2x on top of fusion. These results suggest that the RNG-GEMM overlap methodology is widely applicable towards multiple workload sizes, and also adapts easily to future generations of workloads.
@@ 4.1. Performance Analysis
The performance benefits change with sequence length and embedding dimension sizes (number of heads * head dimension). To understand these trends, we examined the runtime dependencies of each kernel on these parameters:
@@
(a) Individual kernel runtime with different Embedding Dimension: GEMM scales quadratically and Attention/RNG scale linearly.
(b) Individual kernel runtime with different Sequence Length: GEMM scales linearly and Attention/RNG scale quadratically.
Figure 6. Kernel runtime variation with different sequence length and embedding dimension, measured on silicon.
Figure 6 presents the runtime variation for each kernel across different sequence lengths and number of heads. The data shows that while GEMM runtime scales quadratically with the embedding dimension, the Attention and RNG runtime scales linearly. Conversely, Attention and RNG runtime scales quadratically with sequence length, whereas GEMM runtime scales linearly.
With the overall trends of scaling, we now analyze the numbers shown in Figure 5. For RNG-Attention fusion, the absolute runtime saved stays the same regardless of the workload configurations, because Attention and RNG’s complexity increases at the same scale. Since we evaluate the runtime for an entire Transformer Block, fusion brings more saving as sequence length increases which also increases the runtime ratio of Attention/RNG. On the other hand, fusion brings less saving as embedding dimension increases where GEMM runtime scales faster than Attention/RNG runtime. Overall, the RNG-Attention fusion speedup is relatively low, which matches our analysis that they have conflicting hardware resource bottlenecks.
For RNG-GEMM overlap, speedup is correlated with the ratio of sequence length to the embedding dimension, given that the batch size and the hidden dimension don’t change (typically they wouldn’t). The greatest improvement occurs within region 2 in Figure 5, which is the range that common LLM training workloads fall inside.
Figure 7. Analysis of RNG-GEMM overlapping speedup across three regions in Figure 5. Typical LLM training workloads fall inside region 2 with best performance benefits.
Figure 7 illustrates the reasons behind the performance differences of the three regions:
@@
  • Region 2 (optimal speedup): Balances sequence length and embedding dimension, where RNG runtime is shorter but close to the GEMM runtime. Maximum benefits from hiding RNG latency by overlapping with GEMM.
@@
  • Region 3 (decreasing speedup): RNG runtime exceeds GEMM because of the long sequence lengths, leading to full exposure of RNG operations post-GEMM completion.
As newer generations of LLM training workloads typically grow in both input sequence length and embedding dimension, future, bigger workloads are very likely to stay in region 2 where the RNG-GEMM overlapping technique brings high performance benefits. Skewed workloads that fall in region 1 and 3 are much less likely to appear, and even if that happens, the overlapping technique still outperforms the state-of-the-art RNG-Attention fusion.
5. Generalization of Technique
In this section, we explore the implications of the RNG-GEMM overlapping technique with different software and hardware configurations. Specifically, we discuss its performance on future generations of GPUs with higher compute capabilities, and on different implementations of the dropout operator.
5.1. Modeling Future Hardware
We use our analytical model to explore how variations in hardware design will influence the efficiency of the overlapping strategy.
A key aspect of GPU hardware evolution is the consistent increase of computational efficiency with each new generation. For example, the NVIDIA Blackwell GPUs (NVIDIA Corporation, 2024) offer 2x the compute capability as the Hopper generation (NVIDIA Corporation, 2022), but the non-Tensor resources (such as register files, issue stage and execution queues) do not see significant improvements. This trend motivates us to model the potential implications of the RNG-GEMM overlapping technique: the conflicting bottlenecks of RNG and Attention will become a more severe issue, which makes the overlapping approach more favorable.
Figure 8. RNG-GEMM overlap speedup on a Blackwell B100 GPU model with 2x compute capability (TFLOPs). While GEMM runtime decreases, RNG and Attention runtimes remain constant, leading to a greater proportional speedup.
Figure 9. Overlap speedup with increased GPU compute capability (TFLOPs). While GEMM runtime decreases, RNG and Attention runtimes remain constant, leading to a greater proportional speedup.
We evaluated the theoretical performance of our overlapping strategy on a B100 GPU, which offers twice the compute capability of the H100, but other non-Tensor related limiters like the issue pipeline and the ALU pipeline remain unchanged. Computation remains the bottleneck for matrix multiplications, with memory improvements keeping pace with MMA improvements through reduced precision or increased memory bandwidth.
Figure 8 shows our results obtained from modeling workloads with various sequence lengths and embedding dimensions on the B100 chip. Figure 9 illustrates the implications of B100’s increased compute capability on our overlapping technique. Since the most significant runtime reduction is observed in the GEMM kernel, this shift makes RNG latency an even more critical bottleneck, increasing its impact on end-to-end network performance. Although the absolute runtime difference between baseline and overlapping remains similar, the relative speedup improves. This difference can be better observed in workloads with shorter sequence lengths; for longer sequences, where RNG and Attention dominate network runtime, overlapping a shorter GEMM could exacerbate problems by fully exposing RNG latency once GEMM computation completes.
This analysis emphasizes the need for next-generation hardware to consider optimizing traditionally non-Tensor related factors. On the other hand, if hardware optimization on these factors are not possible, we call for future hardware-software co-design to implement kernel overlapping: efficient libraries can be developed to facilitate the overlapping of non-GEMM operations, which typically do not consume extensive computation and memory resources, with GEMM operations.
Figure 10. Stand-alone RNG kernel runtime with different Philox implementations for a sequence length of 16K, measured in silicon. Less iterations lead to shorter runtime.
5.2. Implication of Cheaper RNG
@@
In our prior discussions, we focused on the use of Philox 7 for RNG implementation. This subsection explores the implications of adopting more cost-effective RNG implementations, namely Philox 5 and Philox 3, which have fewer computational iterations and shorter runtimes.
@@
We implemented all three Philox variants (Philox 3, 5, 7) on silicon using the GH100 with the same experimental setup as discussed. We show representative results collected for sequence length = 16K in Figure 10, which has a consistent trend with other configurations not shown on the graph.
@@
Our analytical performance model was validated against the silicon results on overlapping RNG with QKV_GEMM, showing an error margin consistent with previous validations. Using the validated model, we further analyze the implications of varying RNG complexities on overlapping RNG with all four GEMM kernels in the Transformer Block.
Figure 11. Cheaper RNG implementation results in smaller overall speedup.
Figure 12. Overlap speedup with different RNG Philox implementations, derived from theoretical model.
Figure 13. Validation of our analytical model with silicon measurements, showing relative speedups of overlap and fusion methodologies. Data above the 0% horizontal line indicates overlap’s speedup over fusion, and below the line indicates fusion’s speedup over overlap.
As the complexity of the RNG algorithm decreases, the potential for speedup through overlapping also diminishes because the runtimes of the Attention and GEMM kernels remain unchanged. Thus, cheaper RNG implementations should result in smaller overall savings, as shown in Figure 11. This is the case in most scenarios, as shown in Figure 12. However, in certain cases (such as Philox 7 implementation with embedding dimension = 6K and sequence length = 16K) the RNG runtime exceeds the GEMM runtime, introducing performance loss from fully exposed RNG after GEMM completes. This typically occurs with a smaller embedding dimension and a relatively long sequence length - Region 3 as discussed in Figure 5.
@@
6. Validation with Silicon Implementation
We validated the theoretical model’s accuracy with a CUDA implementation on a NVIDIA H100 GPU (NVIDIA Corporation, 2022), specifically using the GH100 HBM3 80GB variant. This GPU supports the latest High Bandwidth Memory (HBM) and provides ample compute resources and memory capacity for LLM training. Across multiple sequence length and embedding dimension configurations, we observed only 2% average difference and 0.04 standard deviation between the analytical model and silicon measurements. The difference is calculated by averaging the absolute difference of speedup between silicon and theoretical results.
6.1. Implementation Details
Our silicon implementation uses production-ready, highly optimized GEMM, Attention, and dropout kernels. The dropout kernel was modified to run in two scenarios: 1) fused within the Attention kernel, and 2) as a stand-alone RNG kernel storing bits representing random numbers in HBM for later use by the Attention kernel. We used the Philox 7 algorithm for RNG.
In terms of kernel scheduling, RNG and GEMM can either run as separate kernels or within the same kernel using warp specialization (Li et al., 2023)(NVIDIA, 2025b)(Kerr et al., 2017). We opted for separate kernels to maintain a clear distinction between independent components in the network such as RNG and GEMM in this case. This approach allows other overlapping strategies to be implemented without deep knowledge of the kernels’ internal complexities, especially for the highly optimized GEMM kernels. Warp specialization enables a more fine-grained overlapping and we anticipate it to bring further, but limited, performance benefits.
The baseline implementation in our experiments includes a QKV_GEMM kernel, followed by a RNG kernel and an Attention kernel (including dropout), all run sequentially.
The fusion implementation includes a stand-alone kernel for QKV_GEMM, and a kernel with fused dropout (including RNG) and Attention, running sequentially on the same CUDA stream. Our optimized kernel fusion minimizes synchronization overhead and maximizes the overlap of RNG operations with Attention’s floating-point computations.
The GPU is warmed up using multiple instances of the same kernel sequence before actual measurements are taken. We validated our CUDA implementation against CPU generated result, ensuring correctness across various batch sizes, sequence lengths, and embedding dimensions.
6.2. Validation Results
Figure 13 shows the comparison of our analytical model and the silicon results. For better illustration purposes, instead of absolute runtime we show relative performance differences between RNG-Attention fusion and RNG-GEMM overlapping to validate our model’s effectiveness for choosing the right technique for each configuration. The analytical model correctly derives the trade-off between the two options. Note that the results obtained here is different from what was shown in Figure 5, because we only overlap with one GEMM instead of four GEMMs in our silicon implementation, simply for validation purposes.
To further validate our assumptions about kernel interference and overhead performance impacts, we conducted the following tests:
  • GEMM Resource Allocation: Typically, the GEMM kernel utilizes all available Register Files and Shared Memory in each SM. Since RNG doesn’t require much resources, in our implementation we carved out 6% of the Registers and 7% of the Shared Memory for the RNG kernel, hypothesizing this would not adversely affect GEMM performance. Our silicon measurements confirmed only 0.5% average performance difference across various GEMM workload sizes.
7. Discussion
7.1. Multi-GPU System
We now extend our previous analysis on single GPU to a multi-GPU system, which is widely used in LLM training since the workload size is huge. The discussed insights for RNG-GEMM overlapping still hold true.
Common strategies for running the workloads in parallel on multi-GPU include Tensor Parallelism (Shoeybi et al., 2019) and Sequence Parallelism (Korthikanti et al., 2023), which split the head and sequence length dimensions, respectively. As depicted in Figure 2, the multi-GPU setting introduces communication layers within each Transformer Block. The communication layers do not degradate the RNG-GEMM overlapping benefits: the latency of these communication layers can be optimized away by overlapping them with the GEMM layers’ computation. Since communication layers utilize distinctly separate resources from GEMM (such as NVLink bandwidth), their concurrent execution will not introduce resource contention and will effectively hide communication latency.
It is also worth noting that parallelism does not change the performance benefit ratios from overlapping RNG and GEMM analyzed earlier. Since the workload is split evenly on each GPU, the ratios of RNG, GEMM and Attention kernel runtime stay the same.
Figure 14. Total memory capacity requirements for stand-alone RNG across different workload configurations.
7.2. Memory Capacity Requirement
Our overlapping methodology requires executing RNG as a stand-alone kernel which stores the generated RNG bits in memory to be later used by the Attention kernel. This is the same amount of memory usage as the baseline, sequential kernel implementation. For each element in Attention’s intermediate matrix, assuming RNG generates 1 bit per element to indicate the dropping status, the entire RNG kernel requires storage of $B*nH*SQ^{2}$ in memory. Figure 14 illustrates the total memory capacity requirements for stand-alone RNG.
Figure 15. Pipelining kernels that reduce memory storage requirements to negligible amount.
Since typical deployment scenario for LLM training involves multiple GPUs, the workload is divided into smaller segments distributed across the GPUs. This division significantly reduces the HBM capacity needed for RNG on each GPU. Figure 14 demonstrates the reduced HBM requirements with parallelism mechanisms. For the mainstream LLM training networks, the required HBM capacity is decreased by a factor of the number of GPUs used, depending on the chosen parallelism strategy and dimension (16 GPUs as shown in the figure).
To further alleviate memory storage concerns, an additional strategy can be applied which involves pipelining the RNG, GEMM, and Attention kernels. Figure 15 illustrates this strategy. It involves processing only a portion of the total computation per kernel at a time, splitting along the sequence length dimension to avoid creating dependencies in the GEMM kernel. With this methodology, we can reduce the RNG-GEMM overlapping’s memory capacity requirements to arbitrarily small segments.
8. Related Work